# KYUSHU-HET-290

### v1

@@ Abstract
Employing the modified Villain lattice formulation of the axion QED, we present an alternative and much simpler derivation of the conclusion of Ref. Honda:2024sdz that the sweep of the axial $U(1)$ non-invertible symmetry operator over the (non-genuine) gauge invariant ’t Hooft line operator with an integer magnetic charge does not leave any effect. The point is that such a ’t Hooft line can be represented by a boundary of a (non-topological) defect that is invariant under the axial transformation on the axion field.
@@ 1 Introduction and summary
The study in Refs. Choi:2022jqy ; Cordova:2022ieu pointed out the possibility that the axial $U(1)$ rotation of the fermion field with fractional angles, although it suffers from the ordinary axial $U(1)$ anomaly, may be understood as a generalized form of the symmetry Gaiotto:2014kfa (see Refs. Schafer-Nameki:2023jdn ; Bhardwaj:2023kri ; Shao:2023gho for reviews), the non-invertible symmetry Aasen:2016dop ; Bhardwaj:2017xup ; Chang:2018iay ; Thorngren:2019iar ; Komargodski:2020mxz ; Koide:2021zxj ; Choi:2021kmx ; Kaidi:2021xfk ; Hayashi:2022fkw ; Choi:2022zal ; Kaidi:2022uux ; Roumpedakis:2022aik ; Bhardwaj:2022yxj ; Cordova:2022ieu ; Choi:2022jqy ; Bhardwaj:2022lsg ; Karasik:2022kkq ; GarciaEtxebarria:2022jky ; Choi:2022fgx ; Yokokura:2022alv ; Nagoya:2023zky ; Anber:2023mlc . In Ref. Honda:2024sdz , the present authors studied how the symmetry operator of the axial $U(1)$ non-invertible symmetry acts on the (non-genuine) ’t Hooft line operator by modeling the axial anomaly by the axion and employing the modified Villain lattice formulation of $U(1)$ gauge theory Sulejmanpasic:2019ytl (see also Ref. Gorantla:2021svj ), which allows a straightforward introduction of the dual $U(1)$ gauge field. In Ref. Honda:2024sdz , it was concluded that the sweep of the symmetry operator over the ’t Hooft line operator does not leave any effect. This conclusion appears inequivalent with the phenomenon concluded in Refs. Choi:2022jqy ; Cordova:2022ieu that the sweep leaves non-trivial traces. Although we believe that under assumptions made in Ref. Honda:2024sdz , our conclusion is correct, the argument in Ref. Honda:2024sdz might be somewhat complicated and, in particular, it appears that the conclusion depends crucially on a miraculous cancellation between the effect of the dressing factor in the ’t Hooft line and the would-be Witten effect Witten:1979ey arising from the axion topological coupling.
@@
In this paper, we present an alternative and much simpler derivation of the conclusion in Ref. Honda:2024sdz , although the basic assumptions are common. Considering the basic importance of this issue in possible applications of the axial $U(1)$ non-invertible symmetry, especially in the form of the selection rules, we believe that it is useful to do this. In this paper, we show that the conclusion in Ref. Honda:2024sdz can be derived from the following observation: The (non-genuine) ’t Hooft line operator considered in Ref. Honda:2024sdz can be realized by a background 2-cocycle gauge field, which couples to the (non-conserved) electric 2-cocycle current $\star f$. Since this coupling is not influenced by the axial transformation on the axion field, $\phi\to\phi+\alpha$, the symmetry operator does not cause any effect on the ’t Hooft line. This simple observation explains the above miraculous cancellation. Also, we hope that the present simple argument gives an idea that our conclusion, although it is shown by employing a particular lattice action, is quite insensitive on the details of the lattice action one adopts. The extension of the present argument to the case of the action of the 1-form non-invertible symmetry in the axion QED Choi:2022fgx ; Yokokura:2022alv , the action of the 1-form non-invertible symmetry operator on the axion string, would be an interesting problem. [^1] Since the (non-genuine) gauge invariant axion string can be represented by a background 1-cochain gauge field introduced in the kinetic term of the axion as we will see in Sect. 2.5, and the 1-form non-invertible symmetry is associated with the shift of the gauge potential, repeating a similar argument as that in the present paper, we expect that the sweep of the symmetry operator over the axion string does not leave any effect.
@@ 2.1 Lattice action
We take the lattice action of Ref. Honda:2024sdz (Eq. (2.24) there) as an explicit starting point for our argument. It is based on the modified Villain formulation of the $U(1)$ lattice gauge theory Sulejmanpasic:2019ytl . After the integration by parts on the periodic hypercubic lattice $\Gamma$, the lattice action can be written as [^2] We denote the hypercube, the cube, the plaquette, and the link on $\Gamma$ by $h$, $c$, $p$, and $l$, respectively. The $p$-cochain, the cup product $\cup$, and the coboundary operator $\delta$ on $\Gamma$ are, roughly speaking, lattice analogues of the $p$-form, the wedge product $\wedge$, and the exterior derivative $\mathrm{d}$ on $T^{4}$, respectively. For precise definitions of the $p$-cochain, the cup product, the higher cup product $\cup_{1}$, the coboundary operator and the “Hodge dual” $\star$ on the hypercubic lattice, see Refs. Chen:2021ppt ; Jacobson:2023cmr . On the cup product, the coboundary operator $\delta$ satisfies the Leibniz rule, $\delta(\alpha\cup\beta)=\delta\alpha\cup\beta+(-1)^{p}\alpha\cup\delta\beta$, where $\alpha$ is a $p$-cochain. The cup product is not commutative on the cochain level but the commutator is expressed by the higher cup as $(-1)^{pq}\beta\cup\alpha=\alpha\cup\beta+(-1)^{p+q}[\delta(\alpha\cup_{1}\beta)-\delta\alpha\cup_{1}\beta-(-1)^{p}\alpha\cup_{1}\delta\beta]$, where $\beta$ is a $q$-cochain.
$S=\sum_{h\in\Gamma}\left\{\frac{1}{g_{0}^{2}}f\cup\star f+i\tilde{a}\cup\delta z+\frac{\mu^{2}}{2}\partial\phi\cup\star\partial\phi+i\delta\ell\cup\chi-i\frac{\mathrm{e}^{2}}{8\pi^{2}}\left[\partial\phi\cup(\mathrm{CS})-4\pi^{2}\phi\cup P_{2}(z)\right]\right\}.$ (2.1)
In this expression,
@@
where $\phi\in\mathbb{R}$ is a 0-cochain and $\ell\in\mathbb{Z}$ is an integer 1-cochain field. $\chi\in\mathbb{R}$ in Eq. (2.1) is another 2-cochain auxiliary field which imposes the another Bianchi identity $\delta\ell=0$. The lattice $U(1)$ Chern–Simons form $(\mathrm{CS})$ is defined by Jacobson:2023cmr  [^3] The motivation for this definition may be understood as follows: In the present lattice formulation, $f\cup f$ is not a total divergence. One can make this total divergence up to $8\pi^{2}\mathbb{Z}$ by adding terms containing $\delta z$ as $f\cup f-4\pi a\cup\delta z+2\pi f\cup_{1}\delta z=\delta(\mathrm{CS})+4\pi^{2}P_{2}(z).$ (2.4) That is, under the Bianchi identity $\delta z=0$, the $\theta$ term defined by $(\theta/8\pi^{2})f\cup f=(\theta/8\pi^{2})[\delta(\mathrm{CS})+4\pi^{2}P_{2}(z)]$ is manifestly topological and changes by $2\pi\mathbb{Z}$ under $\theta\to\theta+2\pi$.
@@
$S^{B}$ $:=\sum_{h\in\Gamma}\biggl{\{}\frac{1}{g_{0}^{2}}\left(f-2\pi B\right)\cup\star\left(f-2\pi B\right)+i\tilde{a}\cup\delta z+\frac{\mu^{2}}{2}\partial\phi\cup\star\partial\phi+i\delta\ell\cup\chi$
@@
The Poincaré dual of $B$ defines a 2-dimensional surface (i.e., defect) $\mathcal{R}$. If $\delta B\neq 0$, the surface $\mathcal{R}$ has the boundary $\gamma$, $\partial\mathcal{R}=\gamma$, along which $\delta B\neq 0$. Then the gauge coupling (2.15) defines a ’t Hooft line operator along $\gamma$. Having these facts in mind, we set
@@
In the present axion QED, $\star f$ is not conserved under the equation of motion, $\delta\star f\neq 0$, and thus the partition function is not invariant under the 1-form gauge transformation on $B$, $B\to B+\delta\omega$. Correspondingly, the Poincarè dual of $B$, the surface $\mathcal{R}$ is not topological, i.e., the partition function with the Boltzmann weight $e^{-S^{B}}$ depends on a precise shape of $\mathcal{R}$. Nevertheless, we may still introduce such a defect and define a (non-genuine) ’t Hooft line operator. Note that, since the combination $(f-2\pi B)\cup\star(f-2\pi B)$ is manifestly invariant under the gauge transformations in Eqs. (2.7)–(2.9), the above method defines an object invariant under those gauge transformations. [^5] The continuum counterpart of Eq. (2.1) possesses an electric $\mathbb{Z}_{\mathrm{e}^{2}}$ 0-form symmetry Hidaka:2020iaz ; Hidaka:2020izy and one can consider a genuine ’t Hooft line operator with a fractional magnetic charge as the boundary of the corresponding symmetry operator. Such a fractionally-charged magnetic monopole worldline cannot be considered in the present lattice formulation with the integer $z$. One would have to consider the “excision method” Abe:2023uan to incorporate such a fractionally-charged ’t Hooft line; the excision method has not completely been developed for 4D gauge theory.
@@ 2.3 Axial $U(1)$ Ward–Takahashi identity
Next, we define the expectation value in the sub-sector by [^11] We assume that, divided by the gauge volumes, the integration measure is defined by $\int\mathrm{D}[\tilde{a}]\mathrm{D}[\phi]\mathrm{D}[\ell]\mathrm{D}[\chi]:=\prod_{l\in\Gamma}\left[\int_{-\pi}^{\pi}\mathrm{d}\tilde{a}(l)\right]\prod_{s\in\Gamma}\left[\int_{-\pi}^{\pi}\mathrm{d}\phi(s)\right]\prod_{l\in\Gamma}\left[\sum_{\ell(l)=-\infty}^{\infty}\right]\prod_{p\in\Gamma}\left[\int_{-\pi}^{\pi}\mathrm{d}\chi(p)\right].$ (2.20)
@@
$\phi\to\begin{cases}\phi+\alpha&\text{for $x\in\mathcal{V}_{4}$},\\ \phi&\text{otherwise}.\\ \end{cases}$ (2.22)
@@ 2.4 Final steps
$\left\langle\exp\left\{\frac{i}{2}\alpha\sum_{c\in\mathcal{M}_{3}}\left[\star j_{5}-\frac{\mathrm{e}^{2}}{4\pi^{2}}(\mathrm{CS})(z)\right]\right\}\right\rangle_{\mathrm{s}}^{B}.$
@@
where $\mathcal{Z}_{\mathcal{M}_{3}}[z]$ is the partition function of a lattice 3D TQFT. In Refs. Honda:2024yte ; Honda:2024sdz , an explicit form of $\mathcal{Z}_{\mathcal{M}_{3}}[z]$ on the lattice is given by a level $N$ BF theory. The basic property of $\mathcal{Z}_{\mathcal{M}_{3}}[z]$, for the configuration (2.26) is, for $\delta z=0$ Honda:2024yte ,
@@
where the full expectation value is defined by [^14] We understand $\int\mathrm{D}[a]\mathrm{D}[z]:=\prod_{l\in\Gamma}\left[\int_{-\pi}^{\pi}\mathrm{d}a(l)\right]\prod_{p\in\Gamma}\left[\sum_{z(p)=-\infty}^{\infty}\right].$ (2.34)
@@
Let us summarize the essential elements in our argument. First, we assumed that the non-genuine but gauge invariant ’t Hooft line operator with an integer magnetic charge can be realized by the defect $B$ (2.16) introduced in the gauge field kinetic term. Next, we assume that the 4D topological part in the axial anomaly $P_{2}(z)$ with the fractional coefficient can be represented by a 3D TQFT as in Eq. (2.31); this enables one to define the symmetry operator as Eq. (2.30). Then, we have the relation (2.33). As far as these assumptions are fulfilled, the above conclusion follows almost independently of the details of the lattice action (i.e., possible modifications of $j_{5}$, $(\mathrm{CS})$, $P_{2}(z)$, and the gauge transformation laws).

### v2

@@ Abstract
Employing the modified Villain lattice formulation of the axion quantum electrodynamics, we present an alternative and much simpler derivation of the conclusion of Ref. Honda:2024sdz that the sweep of the axial $U(1)$ non-invertible symmetry operator over the (non-genuine) gauge invariant ’t Hooft line operator with an integer magnetic charge does not leave any effect. The point is that such a ’t Hooft line can be represented by a boundary of a (non-topological) defect that is invariant under the axial transformation on the axion field.
@@ 1 Introduction and summary
The study in Refs. Choi:2022jqy ; Cordova:2022ieu pointed out the possibility that the axial $U(1)$ rotation of the fermion field with fractional angles, although it suffers from the ordinary axial $U(1)$ anomaly, may be understood as a generalized form of the symmetry Gaiotto:2014kfa (see Refs. Schafer-Nameki:2023jdn ; Bhardwaj:2023kri ; Shao:2023gho for reviews), the non-invertible symmetry Aasen:2016dop ; Bhardwaj:2017xup ; Chang:2018iay ; Thorngren:2019iar ; Komargodski:2020mxz ; Koide:2021zxj ; Choi:2021kmx ; Kaidi:2021xfk ; Hayashi:2022fkw ; Choi:2022zal ; Kaidi:2022uux ; Roumpedakis:2022aik ; Bhardwaj:2022yxj ; Cordova:2022ieu ; Choi:2022jqy ; Bhardwaj:2022lsg ; Karasik:2022kkq ; GarciaEtxebarria:2022jky ; Choi:2022fgx ; Yokokura:2022alv ; Nagoya:2023zky ; Anber:2023mlc . In Ref. Honda:2024sdz , the present authors studied how the symmetry operator of the axial $U(1)$ non-invertible symmetry acts on the (non-genuine) ’t Hooft line operator by modeling the axial anomaly by the axion and employing the modified Villain lattice formulation of $U(1)$ gauge theory Sulejmanpasic:2019ytl (see also Ref. Gorantla:2021svj ), which allows a straightforward introduction of the dual $U(1)$ gauge field. In Ref. Honda:2024sdz , it was concluded that the sweep of the symmetry operator over the ’t Hooft line operator does not leave any effect. This conclusion appears inequivalent with the phenomenon concluded in Refs. Choi:2022jqy ; Cordova:2022ieu that the sweep leaves non-trivial traces. Although we believe that under assumptions made in Ref. Honda:2024sdz , our conclusion is correct, the argument in Ref. Honda:2024sdz might be somewhat complicated and, in particular, it might appear that the conclusion depends crucially on a miraculous cancellation between the effect of the dressing factor in the ’t Hooft line and the would-be Witten effect Witten:1979ey arising from the axion topological coupling.
@@
In this paper, we present an alternative and much simpler derivation of the conclusion in Ref. Honda:2024sdz , although the basic assumptions are common to both. Considering the basic importance of this issue in possible applications of the axial $U(1)$ non-invertible symmetry, especially in the form of the selection rules, we believe that it is useful to do this. In this paper, we show that the conclusion in Ref. Honda:2024sdz can be derived from the following observation: The (non-genuine) ’t Hooft line operator considered in Ref. Honda:2024sdz can be realized by a background 2-cocycle gauge field, which couples to the (non-conserved) electric 2-cocycle current $\star f$. Since this coupling is not influenced by the axial transformation on the axion field, $\phi\to\phi+\alpha$, the symmetry operator does not cause any effect on the ’t Hooft line. This simple observation explains the above miraculous cancellation. Also, we hope that the present simple argument gives an idea that our conclusion, although it is shown by employing a particular lattice action, is quite insensitive on the details of the lattice action one adopts. The extension of the present argument to the case of the action of the 1-form non-invertible symmetry in the axion quantum electrodynamics (QED) Choi:2022fgx ; Yokokura:2022alv , the action of the 1-form non-invertible symmetry operator on the axion string, would be an interesting problem. [^1] Since the (non-genuine) gauge invariant axion string can be represented by a background 1-cochain gauge field introduced in the kinetic term of the axion as we will see in Sect. 2.5, and the 1-form non-invertible symmetry is associated with the shift of the gauge potential, repeating a similar argument as that in the present paper, we expect that the sweep of the symmetry operator over the axion string does not leave any effect.
@@ 2.1 Lattice action
We take the lattice action of Ref. Honda:2024sdz (Eq. (2.24) there) as an explicit starting point for our argument. It is based on the modified Villain formulation of the $U(1)$ lattice gauge theory Sulejmanpasic:2019ytl . After the integration by parts on the periodic hypercubic lattice $\Gamma$, the lattice action can be written as: [^2] We denote the hypercube, the cube, the plaquette, and the link on $\Gamma$ by $h$, $c$, $p$, and $l$, respectively. The $p$-cochain, the cup product $\cup$, and the coboundary operator $\delta$ on $\Gamma$ are, roughly speaking, lattice analogues of the $p$-form, the wedge product $\wedge$, and the exterior derivative $\mathrm{d}$ on $T^{4}$, respectively. For precise definitions of the $p$-cochain, the cup product, the higher cup product $\cup_{1}$, the coboundary operator and the “Hodge dual” $\star$ on the hypercubic lattice, see Refs. Chen:2021ppt ; Jacobson:2023cmr . On the cup product, the coboundary operator $\delta$ satisfies the Leibniz rule, $\delta(\alpha\cup\beta)=\delta\alpha\cup\beta+(-1)^{p}\alpha\cup\delta\beta$, where $\alpha$ is a $p$-cochain. The cup product is not commutative on the cochain level but the commutator is expressed by the higher cup as $(-1)^{pq}\beta\cup\alpha=\alpha\cup\beta+(-1)^{p+q}[\delta(\alpha\cup_{1}\beta)-\delta\alpha\cup_{1}\beta-(-1)^{p}\alpha\cup_{1}\delta\beta]$, where $\beta$ is a $q$-cochain.
$S$ $=\sum_{h\in\Gamma}\biggl{\{}\frac{1}{2g_{0}^{2}}f\cup\star f+i\tilde{a}\cup\delta z+\frac{\mu^{2}}{2}\partial\phi\cup\star\partial\phi+i\delta\ell\cup\chi$
where we assume that the electric $U(1)$ charge $\mathrm{e}$ is an even integer as Ref. Honda:2024sdz . In this expression,
@@
where $\phi\in\mathbb{R}$ is a 0-cochain and $\ell\in\mathbb{Z}$ is an integer 1-cochain field. $\chi\in\mathbb{R}$ in Eq. (2.1) is another 2-cochain auxiliary field which imposes the another Bianchi identity $\delta\ell=0$. The lattice $U(1)$ Chern–Simons form $(\mathrm{CS})$ is defined by Jacobson:2023cmr : [^3] The motivation for this definition may be understood as follows: In the present lattice formulation, $f\cup f$ is not a total divergence. One can make this total divergence up to $8\pi^{2}\mathbb{Z}$ by adding terms containing $\delta z$ as $f\cup f-4\pi a\cup\delta z+2\pi f\cup_{1}\delta z=\delta(\mathrm{CS})+4\pi^{2}P_{2}(z).$ (2.4)
@@
$S^{B}$ $:=\sum_{h\in\Gamma}\biggl{\{}\frac{1}{2g_{0}^{2}}\left(f-2\pi B\right)\cup\star\left(f-2\pi B\right)+i\tilde{a}\cup\delta z+\frac{\mu^{2}}{2}\partial\phi\cup\star\partial\phi+i\delta\ell\cup\chi$
@@
The Poincaré dual of $B$ defines a 2-dimensional surface (i.e. defect) $\mathcal{R}$. If $\delta B\neq 0$, the surface $\mathcal{R}$ has the boundary $\gamma$, $\partial\mathcal{R}=\gamma$, along which $\delta B\neq 0$. Then the gauge coupling (2.15) defines a ’t Hooft line operator along $\gamma$. Having these facts in mind, we set
@@
In the present axion QED, $\star f$ is not conserved under the equation of motion, $\delta\star f\neq 0$, and thus the partition function is not invariant under the 1-form gauge transformation on $B$, $B\to B+\delta\omega$. Correspondingly, the Poincarè dual of $B$, the surface $\mathcal{R}$ is not topological, i.e. the partition function with the Boltzmann weight $e^{-S^{B}}$ depends on a precise shape of $\mathcal{R}$. Nevertheless, we may still introduce such a defect and define a (non-genuine) ’t Hooft line operator. Note that, since the combination $(f-2\pi B)\cup\star(f-2\pi B)$ is manifestly invariant under the gauge transformations in Eqs. (2.7)–(2.9), the above method defines an object invariant under those gauge transformations. [^5] The continuum counterpart of Eq. (2.1) possesses an electric $\mathbb{Z}_{\mathrm{e}^{2}}$ 1-form (invertible) symmetry Hidaka:2020iaz ; Hidaka:2020izy and one can consider a ’t Hooft line operator with a fractional magnetic charge as the boundary of the corresponding symmetry operator. Such a fractionally-charged magnetic monopole worldline cannot be considered in the present lattice formulation with the integer $z$. One would have to consider the “excision method” Abe:2023uan to incorporate such a fractionally-charged ’t Hooft line; the excision method has not completely been developed for 4D gauge theory.
@@ 2.3 Axial $U(1)$ Ward–Takahashi identity
Next, we define the expectation value in the sub-sector by: [^11] We assume that, divided by the gauge volumes, the integration measure is defined by $\int\mathrm{D}[\tilde{a}]\mathrm{D}[\phi]\mathrm{D}[\ell]\mathrm{D}[\chi]:=\prod_{l\in\Gamma}\left[\int_{-\pi}^{\pi}\mathrm{d}\tilde{a}(l)\right]\prod_{s\in\Gamma}\left[\int_{-\pi}^{\pi}\mathrm{d}\phi(s)\right]\prod_{l\in\Gamma}\left[\sum_{\ell(l)=-\infty}^{\infty}\right]\prod_{p\in\Gamma}\left[\int_{-\pi}^{\pi}\mathrm{d}\chi(p)\right].$ (2.20)
@@
$\phi(s)\to\begin{cases}\phi(s)+\alpha&\text{for $s\in\mathcal{V}_{4}$},\\ \phi(s)&\text{otherwise}.\\ \end{cases}$ (2.22)
@@ 2.4 Final steps
$\left\langle\exp\left\{\frac{i}{2}\alpha\sum_{c\in\mathcal{M}_{3}}\left[\star j_{5}-\frac{\mathrm{e}^{2}}{4\pi^{2}}(\mathrm{CS})\right]\right\}\right\rangle_{\mathrm{s}}^{B}.$
@@
where $\mathcal{Z}_{\mathcal{M}_{3}}[z]$ is the partition function of a lattice 3D TQFT. In Refs. Honda:2024yte ; Honda:2024sdz , an explicit form of $\mathcal{Z}_{\mathcal{M}_{3}}[z]$ on the lattice is given by a level $N$ BF theory. The basic property of $\mathcal{Z}_{\mathcal{M}_{3}}[z]$, for the configuration in Eq. (2.26) is, for $\delta z=0$ Honda:2024yte ,
@@
where the full expectation value is defined by: [^14] We understand $\int\mathrm{D}[a]\mathrm{D}[z]:=\prod_{l\in\Gamma}\left[\int_{-\pi}^{\pi}\mathrm{d}a(l)\right]\prod_{p\in\Gamma}\left[\sum_{z(p)=-\infty}^{\infty}\right].$ (2.34)
@@
Let us summarize the essential elements in our argument. First, we assumed that the non-genuine but gauge invariant ’t Hooft line operator with an integer magnetic charge can be realized by the defect $B$ (2.16) introduced in the gauge field kinetic term. Next, we assumed that the 4D topological part in the axial anomaly $P_{2}(z)$ with the fractional coefficient can be represented by a 3D TQFT as in Eq. (2.31); this enables one to define the symmetry operator as Eq. (2.30). Then, we have the relation (2.33). As far as these assumptions are fulfilled, the above conclusion follows almost independently of the details of the lattice action (i.e., possible modifications of $j_{5}$, $(\mathrm{CS})$, $P_{2}(z)$, and the gauge transformation laws).
@@ 2.5 Axion string
Summarizing this paper, we have shown the same conclusions as Ref. Honda:2024sdz on the action of the non-invertible symmetry operator in the axion QED on magnetic objects (the ’t Hooft line and the axion string) in a much simpler argument.
Note added
In this paper and in Ref. Honda:2024sdz , we have considered the situation in which one employs only degrees of freedom in the axion QED. The gauge invariance of line and surface operators then forces us to consider non-genuine operators. If we regard the axion QED as a low-energy effective theory of the QED with the Peccei–Quinn scalar, however, there arises a natural possibility that there exist massless degrees of freedom residing only on line and surface operators and they make operators genuine. Then, the action of the non-invertible symmetry operator can be different. We are grateful to Yuya Tanizaki for helpful discussions on these points. In Appendix A, we give an example of a gauge invariant genuine ’t Hooft line operator on the lattice and show that with that definition the sweep of the symmetry operator produces a surface integral of the $U(1)$ field strength as discussed in Refs. Choi:2022jqy ; Cordova:2022ieu ; Choi:2022fgx .